### v1

Abstract
This paper presents almanac and ephemeris message representation for lunar navigation satellites supporting the Lunar Augmented Navigation System (LANS). The proposed method combines osculating orbital elements, Chebyshev polynomials, and Fourier series to efficiently represent lunar satellite trajectories subject to complex perturbations from lunar gravity and third-body effects. For the ephemeris, a hybrid Chebyshev–Fourier formulation improves fitting accuracy over long arcs while maintaining message compactness under the datasize constraint of the LunaNet Interoperability Specification. For the almanac, a compact low-order polynomial and Fourier model is introduced to capture mid-term orbital variations over a 15-day fitting arc. The approach is validated for multiple orbit regimes, including 30-hour, 24-hour, and 12-hour elliptical Lunar frozen orbits (ELFOs) and a 6-hour polar orbit. Results show that the proposed framework achieves sub-meter position and sub-millimeter-per-second velocity fitting errors within the 900-bit limit for 6-hour ephemeris arcs, and almanac fitting accuracy sufficient for reliable satellite visibility identification in warm-start operations.
{IEEEkeywords}
Lunar positioning, navigation, and timing (PNT), almanac, ephemeris, GNSS
1 Introduction
The pace of lunar exploration has accelerated rapidly in recent years, with more than 40 missions planned by space agencies and private entities between 2024 and 2030 [1]. To support these activities, robust and precise positioning, navigation, and timing (PNT) services are indispensable for orbit determination, surface mobility, logistics, and communication across the lunar environment. In response, space agencies are jointly developing LunaNet—an interoperable “network of networks” designed to provide navigation and communication services both in lunar orbit and on the surface [2]. A key component of LunaNet’s navigation services is the Lunar Augmented Navigation Service (LANS), a constellation of lunar-orbiting navigation satellites that will broadcast Augmented Forward Signals (AFS) to enable users to determine their position and time [3]. National Aeronautics and Space Administration (NASA), European Space Agency (ESA), and Japan Aerospace Exploration Agency (JAXA) each plan to deploy their own LANS constellations as part of LunaNet: the Lunar Communications Relay and Navigation Systems (LCRNS) [4], the Lunar Communications and Navigation Service (LCNS) [5], and the Lunar Navigation Satellite System (LNSS) [6], respectively.
[…]
Table 1: Summary of Prior Works on Ephemeris Design for Lunar Navigation Satellites (OE: Osculating Elements)
Reference Ephemeris Almanac Orbits Position Velocity Datasize
[7] Chebyshev ELFO (12 hr), LLO
[11] OE + Empirical Acceleration, Chebyshev ELFO (23.95 hr)
[12] OE + Chebyshev ELFO (24.37 hr)
[9] OE + Chebyshev ELFO (12, 18, 24 hr), DRO
This work OE + Chebyshev + Fourier OE (Polynomial + Fourier) ELFO (12, 24, 30 hr), Polar
Several studies have explored approaches to parameterize lunar navigation satellite orbits. [7] demonstrated a Chebyshev-based state-fitting method that directly fits the Cartesian position and velocity components subject to state constraints, achieving meter-level accuracy for a 12-hour ELFO. Similarly, [11] proposed using Chebyshev polynomials and osculating orbital elements augmented with empirical accelerations to model the orbits of LCNS satellites in 23.95-hour ELFOs. [12] and [9] introduced hybrid ephemeris representations combining osculating elements and Chebyshev polynomials, demonstrating that such formulations can achieve longer fitting durations and smaller data sizes than Chebyshev-only methods. These prior studies are summarized in Table 1.
While these prior studies have established valuable foundations for lunar ephemeris design, their achievable validity durations remain limited—typically two hours for 12-hour ELFOs and four hours for 24-hour ELFOs [9]. Given the scarcity of lunar tracking assets, extending the ephemeris validity (e.g., to four hours or longer) would be beneficial to reduce the operational burden of frequent updates. Furthermore, none of the previous works have examined almanac design for lunar navigation satellites, which is essential for cold- and warm-start receiver operations.
This paper introduces a unified framework for ephemeris and almanac design tailored for LunaNet satellites. The proposed ephemeris representation integrates osculating orbital elements with Chebyshev polynomials and Fourier series to capture both short- and mid-term orbital perturbations, improving fitting accuracy over longer intervals compared to prior methods. For the almanac, a compact polynomial–Fourier model is developed to represent the long-term evolution of osculating elements and argument of latitude, achieving sub-degree median latitude and longitude errors over 15-day arcs—sufficient for warm-start satellite visibility identification. We validate the ephemeris and almanac generation across multiple orbit regimes: 12-, 24-, and 30-hour ELFOs and low-altitude polar orbits.
The remainder of this paper is organized as follows. Section 2 describes the reference frames, dynamics models, and target lunar orbits used in the analysis. Section 3 presents the proposed ephemeris and almanac representations. Section 4 details the fitting methodology and message-size computation. Sections 5 and 6 present the ephemeris and almanac fitting results, respectively. Finally, Section 7 concludes with key findings and directions for future work.
2 Orbit Analysis
2.1 Coordinate Frames and Dynamics
We consider three coordinate frames for the orbit analysis: the Moon-Centered Inertial (MCI) frame, the Moon Principal Axis (PA) frame, and the Principal Axis Inertial (PAI) frame.
The MCI frame is defined with its origin at the Moon’s center of mass, and its axes are aligned with those of the Geocentric Celestial Reference Frame (GCRF) at a reference epoch. This frame serves as the inertial reference for orbit propagation and dynamical modeling.
The PA frame is a rotating frame whose axes are aligned with the Moon’s principal axes of inertia at each epoch $t$. This frame rotates with the Moon and is used as the output frame for the satellite position and velocity in the ephemeris representation [13].
The PAI frame is also centered at the Moon’s center of mass, but its axes are aligned with the instantaneous orientation of the Moon’s principal axes of inertia at each time $t$, without accounting for the rotational motion of the Moon itself. Because the Moon’s principal axes slowly precess over time, a distinct PAI frame exists at each epoch. This quasi-inertial frame is used during the fitting process to generate osculating orbital elements.
The position vectors in the MCI, PAI, and PA frames are denoted as $\mathbf{r}^{\mathrm{MCI}}$, $\mathbf{r}^{\mathrm{PAI}}$, and $\mathbf{r}^{\mathrm{PA}}$, respectively. Similarly, the velocity vectors are denoted as $\mathbf{v}^{\mathrm{MCI}}$, $\mathbf{v}^{\mathrm{PAI}}$, and $\mathbf{v}^{\mathrm{PA}}$. The positions in the PA and PAI frames are identical, but their velocities differ by the cross product of the angular velocity of the PA frame with respect to the inertial frame and the position vector. The transformations between the MCI, PAI, and PA frames are summarized in [13].
The lunar orbit propagation is performed in the MCI frame using a dynamics model that includes the Moon’s gravitational field up to degree and order 80, as well as third-body perturbations from the Earth and the Sun [14].
2.2 Orbit Propagation Analysis
Figure 1: Time histories of osculating orbital elements for the 30-hour ELFO (LCRNS) over 365 days.
Figure 2: Fourier spectra of the osculating orbital elements for the 30-hour ELFO (LCRNS) over 365 days.
Figure 3: Time histories of osculating orbital elements for the 6-hour polar orbit over 365 days.
Figure 4: Fourier spectra of the osculating orbital elements for the 6-hour polar orbit over 365 days.
Fig. 1 presents the time histories of the osculating orbital elements (excluding mean anomaly) for the 30-hour Elliptical Lunar Frozen Orbit (ELFO) corresponding to the LCRNS reference orbit, propagated over 365 days. As designed, the semi-major axis, eccentricity, inclination, and argument of periapsis remain nearly constant throughout the year, confirming the long-term stability of the frozen-orbit configuration. In contrast, the longitude of the ascending node exhibits a clear secular regression, primarily driven by the Moon’s non-spherical gravity field and its rotational coupling effects.
The corresponding Fourier spectra, shown in Fig. 2, reveal dominant periodic components near half the lunar sidereal period ($\approx 13.7$ days) in $e$, $i$, $\Omega$, and $\omega$. These components are consistent with third-body perturbations induced by Earth. In contrast, the semi-major axis spectrum exhibits high-frequency power near the orbital period ($\approx 30$ hours), reflecting short-period variations arising from higher-order terms in the lunar gravity potential.
The results for the low-altitude 6-hour polar orbit are shown in Figs 3 and 4. Similar to the ELFO case, the semi-major axis remains relatively stable, while the eccentricity, inclination, and argument of periapsis exhibit gradual secular variations due to higher-order lunar gravitational harmonics. A strong periodic component near half the lunar sidereal period ($\approx 13.7$ days) persists in the inclination spectrum, again indicative of third-body coupling. However, for $e$ and $\omega$, the dominant spectral power shifts toward the orbital-period band ($\approx 6$ hours), highlighting the increasing influence of short-period terms in low-altitude regimes.
3 Ephemeris and Almanac Representations
3.1 Ephemeris Representation
The proposed ephemeris representation combines osculating orbital elements with Chebyshev polynomials and Fourier series to model the satellite’s position and velocity over a specified time interval. Similar to the ephemeris design presented in [12, 9], we represent the two-body motion using five osculating orbital elements: semi-major axis $a$, eccentricity $e$, inclination $i$, longitude of ascending node $\lambda$, and argument of periapsis $\omega$, while the mean anomaly $M$ and argument of latitude $u$ are computed from these elements and time.
The orbital elements are converted to Cartesian position and velocity in the PA frame. Then, the residuals between the actual ephemeris data and the two-body motion in the PA frame are modeled using Chebyshev polynomials and Fourier series to capture both secular trends and periodic variations.
Chebyshev polynomials of the first kind, denoted $F_{n}(\tau)$, form an orthogonal basis on the interval $\tau\in[-1,1]$, where
$\tau=\frac{t-t_{0}}{T_{\mathrm{fit}}},$ (1)
with $T_{\mathrm{fit}}$ the fitting interval and $t_{0}$ the reference epoch at its midpoint. They are defined recursively as
$F_{0}(\tau)=1,\quad F_{1}(\tau)=\tau$ (2)
$F_{n+1}(\tau)=2\tau F_{n}(\tau)-F_{n-1}(\tau),\quad n\geq 1.$
Chebyshev polynomials minimize the maximum approximation error (the minimax property) [15], allowing high-accuracy fitting with a relatively small number of coefficients. An $n$th-order polynomial requires $n+1$ coefficients to represent the function over the normalized interval.
The second component of the ephemeris representation is the Fourier series, which captures periodic variations in the satellite’s motion, which is given by:
$f(t)=c_{0}+\left[C_{c}\cos\left(2u(t)\right)+C_{sn}\sin\left(2u(t)\right)\right],$ (3)
where $c_{0}$ is the average value of the function over one period, which will be shared with the Chebyshev polynomials, $u(t)$ is the argument of latitude at time $t$, and $C_{cn}$ and $C_{sn}$ are the Fourier coefficients for the cosine and sine terms, respectively. We use the argument of latitude to capture the periodic effects corresponding to the orbital period. The longer periodic effects which has period of the lunar sidereal period can be sufficiently captured by the polynomial terms.
The algorithm to compute the x, y, z position and velocity in the PA frame from the ephemeris parameters is summarized in Algorithm 1.
1
2Compute the mean anomaly at time $t$:
$M(t)=M_{0}+n(t-t_{0}),\quad n=\sqrt{\mu_{\leftmoon}/a_{0}^{3}}.$
3Solve Kepler’s equation for $E(t)$:
M(t)=E(t)esinE(t)M(t)=E(t)-esinE(t)
4Compute the true anomaly:
$\nu(t)=2\arctan\!\left(\sqrt{\frac{1+e_{0}}{1-e_{0}}}\tan\frac{E(t)}{2}\right).$
5Compute the orbital-plane radius, argument of latitude, and position in the plane:
$r(t)=a(t)\big(1-e_{0}\cos E(t)\big),\quad u(t)=w_{0}+\nu(t).$
$x^{p}(t)=r(t)\cos u(t),\quad y^{p}(t)=r(t)\sin u(t).$
6Adjust the longitude of the ascending node for precession:
$\lambda(t)=\lambda_{0}-\omega_{b}(t-t_{0}).$
7Compute the nominal position in the PA frame:
$x^{\mathrm{PA}}_{0}(t)$ $=x^{p}(t)\cos\lambda(t)-y^{p}(t)\cos i_{0}\sin\lambda(t)$
$y^{\mathrm{PA}}_{0}(t)$ $=x^{p}(t)\sin\lambda(t)+y^{p}(t)\cos i_{0}\cos\lambda(t)$
$z^{\mathrm{PA}}_{0}(t)$ $=y^{p}(t)\sin i_{0}$
8Compute the nominal velocity in the PA frame:
$\dot{x}^{\mathrm{PA}}_{0}(t)$ $=-x^{p}(t)\sin\lambda(t)\dot{\lambda}(t)-y^{p}(t)\cos i_{0}\cos\lambda(t)\dot{\lambda}(t)$
$+\dot{x}^{p}(t)\cos\lambda(t)-\dot{y}^{p}(t)\cos i_{0}\sin\lambda(t)$
$\dot{y}^{\mathrm{PA}}_{0}(t)$ $=x^{p}(t)\cos\lambda(t)\dot{\lambda}(t)-y^{p}(t)\cos i_{0}\sin\lambda(t)\dot{\lambda}(t)$
$+\dot{x}^{p}(t)\sin\lambda(t)+\dot{y}^{p}(t)\cos i_{0}\cos\lambda(t)$
$\dot{z}^{\mathrm{PA}}_{0}(t)$ $=\dot{y}^{p}(t)\sin i_{0}$
where
$\dot{\lambda}(t)$ $=-\omega_{b},\quad\ \dot{r}(t)=ae\sin E(t)\dot{E}$
$\dot{u}(t)$ $=\frac{\dot{E}\sqrt{1-e^{2}}}{1-e\cos E(t)},\quad\dot{E}(t)=\frac{n}{1-e\cos E(t)}$
$\dot{x}^{p}(t)$ $=\dot{r}(t)\cos u(t)-r(t)\sin u(t)\dot{u}(t)$
$\dot{y}^{p}(t)$ $=\dot{r}(t)\sin u(t)+r(t)\cos u(t)\dot{u}(t),$
9Compute the residuals in the PA frame $(\xi=\{x,y,z\})$:
$\Delta{\xi}^{\mathrm{PA}}(t)$ $=\sum_{k=0}^{N}C_{F,k}^{\xi}F_{k}\left(\frac{t-t_{0}}{T_{fit}}\right)$
$+C_{c}^{\xi}\cos(u(t))+C_{s}^{\xi}\sin(u(t))$
$\Delta\dot{\xi}^{\mathrm{PA}}(t)$ $=\frac{1}{T_{fit}}\sum_{k=0}^{N}C_{F,k}^{\xi}\dot{F}_{k}\left(\frac{t-t_{0}}{T_{fit}}\right)$
$-C_{c}^{\xi}\sin(u(t))\dot{u}(t)+C_{s}^{\xi}\cos(u(t))\dot{u}(t)$
10Compute the position and velocity in the PA frame $(\xi=\{x,y,z\})$:
$\xi^{\mathrm{PA}}(t)=\xi^{\mathrm{PA}}_{0}(t)+\Delta\xi^{\mathrm{PA}}(t).$
$\dot{\xi}^{\mathrm{PA}}(t)=\dot{\xi}^{\mathrm{PA}}_{0}(t)+\Delta\dot{\xi}^{\mathrm{PA}}(t).$
Algorithm 1 Ephemeris to State Conversion
3.2 Almanac Representation
The almanac provides a simplified, long-term representation of the satellite’s orbit that preserves its essential geometric characteristics while minimizing computational complexity and message size. Unlike the ephemeris, which fits short-arc dynamics using residuals from a two-body model, the almanac fitting is applied directly to the osculating orbital elements over extended time intervals of several days. In this regime, lunar gravitational harmonics and third-body perturbations introduce significant deviations from two-body motion, making direct element fitting more robust and compact.
To capture both secular and periodic variations, each osculating element $\xi(t)$ is expressed as a combination of a low-order polynomial and a single-term Fourier series:
$\xi(t)$ $=\beta_{0}^{\xi}+\beta_{1}^{\xi}t+\left[C_{c}^{\xi}\cos\!\left(\omega t\right)+C_{s}^{\xi}\sin\!\left(\omega t\right)\right],$ (4)
[…]
where $\xi(t)$ denotes the osculating orbital element at time $t$; $\beta_{k}^{\xi}$ are the polynomial coefficients of order $k$; and $C_{c}^{\xi}$ and $C_{s}^{\xi}$ are the Fourier coefficients for the cosine and sine terms, respectively. The frequency $\omega$ is chosen according to the dominant periodicity observed in the Fourier analysis (Section 2): for the semi-major axis, the orbital period $T_{\mathrm{orb}}$ captures short-period variations, whereas for the remaining elements ($e,i,\Omega,\omega,M,u$), the lunar sidereal period $T_{\mathrm{sid}}=27.321661\penalty 10000\ \text{days}$ captures longer-period modulations driven by third-body and gravitational coupling effects.
Each element’s parameter set is defined as
$\bm{\alpha}_{\xi}=[\beta_{0}^{\xi},\beta_{1}^{\xi},C_{c}^{\xi},C_{s}^{\xi}],$
for each osculating element and $u$.
This compact polynomial–Fourier formulation enables accurate prediction of satellite geometry over multi-day periods while maintaining a lightweight parameterization suitable for broadcast almanac messages. The algorithm to compute the Cartesian position and velocity components $(x,y,z,\dot{x},\dot{y},\dot{z})$ in the PA frame from the almanac parameters is provided in Algorithm 2.
1
2Compute the orbital period used for the computation of the semi-major axis:
$T_{orb}(t)=2\pi\sqrt{\frac{\bar{a}(t)^{3}}{\mu_{\leftmoon}}},\qquad\bar{a}(t)=\beta_{0}^{a}+\beta_{1}^{a}.$
3Compute the angular frequencies for the Fourier terms:
$\omega_{orb}=\frac{2\pi}{T_{orb}(t)},\qquad\omega_{sid}=\frac{4\pi}{T_{syn}}.$
4Compute the time-varying orbital elements:
$a(t)$ $=\beta_{0}^{a}+\beta_{1}^{a}t+\big[C_{c}^{a}\cos(\omega_{orb}t)+C_{s}^{a}\sin(\omega_{orb}t)\big]$
$e(t)$ $=\beta_{0}^{e}+\beta_{1}^{e}t+\big[C_{c}^{e}\cos(\omega_{sid}t)+C_{s}^{e}\sin(\omega_{sid}t)\big]$
$i(t)$ $=\beta_{0}^{i}+\beta_{1}^{i}t+\big[C_{c}^{i}\cos(\omega_{sid}t)+C_{s}^{i}\sin(\omega_{sid}t)\big]$
$\omega(t)$ $=\beta_{0}^{\omega}+\beta_{1}^{\omega}t+\big[C_{c}^{\omega}\cos(\omega_{sid}t)+C_{s}^{\omega}\sin(\omega_{sid}t)\big]$
$\lambda(t)$ $=\beta_{0}^{\lambda}+\beta_{1}^{\lambda}t+\big[C_{c}^{\lambda}\cos(\omega_{sid}t)+C_{s}^{\lambda}\sin(\omega_{sid}t)\big]$
$\quad-\omega_{b}t,$
5Compute the nominal mean anomaly using trapezoidal integration:
$n(t)=\sqrt{\frac{\mu_{\leftmoon}}{a(t)^{3}}},\qquad M_{0}(t)=\int_{0}^{t}n(\tau)\,d\tau.$
6Correct the mean anomaly using polynomial and sidereal Fourier terms:
$M(t)$ $=M_{0}(t)+\beta_{0}^{M}+\beta_{1}^{M}t$
$\quad+\big[C_{c}^{M}\cos(\omega_{sid}t)+C_{s}^{M}\sin(\omega_{sid}t)\big].$
7Solve Kepler’s equation for $E(t)$:
M(t)=E(t)esinE(t)M(t)=E(t)-esinE(t)
8Compute the true anomaly:
$\nu(t)=2\arctan\!\left(\sqrt{\frac{1+e(t)}{1-e(t)}}\tan\frac{E(t)}{2}\right).$
9Compute the argument of latitude:
$u(t)$ $=\omega(t)+\nu(t)+\beta_{0}^{u}+\beta_{1}^{u}t$
$\quad+\big[C_{c}^{u}\cos(\omega_{sid}t)+C_{s}^{u}\sin(\omega_{sid}t)\big].$
10Compute the orbital-plane radius and position:
$r(t)$ $=a(t)\big(1-e(t)\cos E(t)\big)$
$x^{p}(t)$ $=r(t)\cos u(t)$
$y^{p}(t)$ $=r(t)\sin u(t)$
11Compute the position in the PA frame:
$x^{\mathrm{PA}}(t)$ $=x^{p}(t)\cos\lambda(t)-y^{p}(t)\cos i(t)\sin\lambda(t)$
$y^{\mathrm{PA}}(t)$ $=x^{p}(t)\sin\lambda(t)+y^{p}(t)\cos i(t)\cos\lambda(t)$
$z^{\mathrm{PA}}(t)$ $=y^{p}(t)\sin i(t)$
12Compute the velocity in the PA frame by finite differencing $(\xi=\{x,y,z\})$:
$\dot{\xi}^{\mathrm{PA}}(t)=\frac{\xi^{\mathrm{PA}}(t+\Delta t)-\xi^{\mathrm{PA}}(t)}{\Delta t}.$
Algorithm 2 Almanac to State Conversion
4 Optimization
4.1 Least Squares Fitting
Both the ephemeris and almanac parameterizations are optimized via a least-squares fitting procedure. Given a set of $N_{T}$ sampled position and velocity states $\{\mathbf{X}^{\mathrm{PAI}}(t_{m})\}_{m=1}^{N_{T}}$ in the Principal-Axis Inertial (PAI) frame over a fitting interval $t_{\mathrm{init}},t_{f}]$, where $\mathbf{X}^{\mathrm{PAI}}(t_{m})=[x^{\mathrm{PAI}}(t_{m}),y^{\mathrm{PAI}}(t_{m}),z^{\mathrm{PAI}}(t_{m})]^{\top}$, the goal is to determine the optimal parameter set $\bm{\alpha}$ that minimizes the residuals between the sampled data and the model prediction.
The sampling epochs are distributed using Chebyshev–Lobatto nodes [16],
$t_{m}=\frac{t_{f}+t_{\mathrm{init}}}{2}+\frac{t_{f}-t_{\mathrm{init}}}{2}\cos\!\left(\frac{m\pi}{N_{T}}\right),$ (5)
to reduce Runge’s phenomenon and improve numerical stability during polynomial fitting. The fitting interval duration is denoted $T_{\mathrm{fit}}=t_{f}-t_{\mathrm{init}}$.
4.1.1 Ephemeris Fitting
The ephemeris fitting procedure is performed as follows:
  1. 1.
    Define the reference epoch $t_{0}=(t_{\mathrm{init}}+t_{f})/2$ and initialize the osculating orbital elements $\{a_{0},e_{0},i_{0},\Omega_{0},\omega_{0},M_{0}\}$ from the PAI state at $t_{0}$.
  2. 2.
    Using Algorithm 1 (lines 1–6), propagate the nominal two-body motion to obtain the predicted state $\mathbf{X}^{\mathrm{PA}}_{0}(t_{m})$ and argument of latitude $u_{m}$ at each node $t_{m}$.
  3. 3.
    Compute the residuals between the sampled and nominal states:
    $\Delta\mathbf{X}^{\mathrm{PAI}}(t_{m})$ $=\mathbf{X}^{\mathrm{PAI}}(t_{m})-\mathbf{X}^{\mathrm{PA}}_{0}(t_{m})$ (6)
    $\quad(m=1,\ldots,N_{T}).$
  4. 4.
    Construct the design matrix $T$ based on the Chebyshev–Fourier basis:
    $T=$ (7)
    $\begin{bmatrix}F_{0}(\tau_{1})&\cdots&F_{n}(\tau_{1})&\cos(2u_{1})&\sin(2u_{1})\\ F_{0}(\tau_{2})&\cdots&F_{n}(\tau_{2})&\cos(2u_{2})&\sin(2u_{2})\\ \vdots&\ddots&\vdots&\vdots&\vdots\\ F_{0}(\tau_{N_{T}})&\cdots&F_{n}(\tau_{N_{T}})&\cos(2u_{N_{T}})&\sin(2u_{N_{T}})\end{bmatrix}$,
    where $\tau_{m}=(t_{m}-t_{0})/T_{\mathrm{fit}}$ and $F_{n}(\tau)$ denotes the $n$th Chebyshev polynomial.
  5. 5.
    Solve the least-squares problem to estimate the ephemeris parameters:
    $\min_{\bm{\alpha}}\sum_{m=1}^{N_{T}}\left\|\Delta\mathbf{X}^{\mathrm{PAI}}(t_{m})-T_{m}\bm{\alpha}\right\|_{2}^{2},$ (8)
    where $T_{m}$ denotes the $m$th row of $T$.
4.1.2 Almanac Fitting
The almanac parameters are obtained through a similar least-squares process applied directly to each osculating orbital element:
  1. 1.
    Compute the osculating elements $\{a(t_{m}),e(t_{m}),i(t_{m}),\Omega(t_{m}),\omega(t_{m}),M(t_{m})\}$ from the PAI states.
  2. 2.
    Define the residuals for each element as
    $\Delta\xi(t_{m})=\begin{cases}\xi(t_{m}),&\xi\in\{a,e,i,\omega\},\\[4.0pt] \xi(t_{m})+\omega_{b}t_{m},&\xi=\Omega,\end{cases}$ (9)
    where $\omega_{b}=2.6617\times 10^{-6}\penalty 10000\ \mathrm{rad/s}$ is the lunar rotation rate.
  3. 3.
    Compute the mean semi-major axis over the fitting interval:
    $\bar{a}=\frac{1}{N_{T}}\sum_{m=1}^{N_{T}}a(t_{m}).$ (10)
  4. 4.
    Construct the design matrix for the semi-major axis:
    $T_{a}$ $=$ (11)
    $\begin{bmatrix}1&t_{1}&\cos(\omega_{\mathrm{orb}}t_{1})&\sin(\omega_{\mathrm{orb}}t_{1})\\ 1&t_{2}&\cos(\omega_{\mathrm{orb}}t_{2})&\sin(\omega_{\mathrm{orb}}t_{2})\\ \vdots&\vdots&\vdots&\vdots\\ 1&t_{N_{T}}&\cos(\omega_{\mathrm{orb}}t_{N_{T}})&\sin(\omega_{\mathrm{orb}}t_{N_{T}})\end{bmatrix}$,
    where $\omega_{\mathrm{orb}}=2\pi/T_{\mathrm{orb}}$ and $T_{\mathrm{orb}}=2\pi\sqrt{\bar{a}^{3}/\mu_{\leftmoon}}$.
  5. 5.
    Solve the least-squares problem for the semi-major axis:
    $\min_{\bm{\alpha}_{a}}\sum_{m=1}^{N_{T}}\|\Delta a(t_{m})-T_{a,m}\bm{\alpha}_{a}\|_{2}^{2}.$ (12)
  6. 6.
    Repeat Steps 3–5 for each remaining element $\xi\in\{e,i,\omega,\Omega\}$ using
    $T_{\xi}$ $=$ (13)
    $\begin{bmatrix}1&t_{1}&\cos(\omega_{\mathrm{sid}}t_{1})&\sin(\omega_{\mathrm{sid}}t_{1})\\ 1&t_{2}&\cos(\omega_{\mathrm{sid}}t_{2})&\sin(\omega_{\mathrm{sid}}t_{2})\\ \vdots&\vdots&\vdots&\vdots\\ 1&t_{N_{T}}&\cos(\omega_{\mathrm{sid}}t_{N_{T}})&\sin(\omega_{\mathrm{sid}}t_{N_{T}})\end{bmatrix}$,
    where $\omega_{\mathrm{sid}}=4\pi/T_{\mathrm{sid}}$ and $T_{\mathrm{sid}}=27.321661\penalty 10000\ \text{days}$ is the lunar sidereal period.
  7. 7.
    Evaluate the mean anomaly $M_{0}(t_{m})$ using Algorithm 2 (lines 1–4) and compute $\Delta M=M(t_{m})-M_{0}(t_{m})$. Fit $\bm{\alpha}_{M}$ using the same design matrix $T_{M}=T_{\xi}$.
  8. 8.
    Finally, compute the argument of latitude residuals $\Delta u=u(t_{m})-[\omega(t_{m})+\nu(t_{m})]$, and solve for $\bm{\alpha}_{u}$ using $T_{u}=T_{\xi}$.
Figure 5: Data size versus 95th-percentile position and velocity errors for the 30-hour ELFO (LCRNS) across fitting lengths and parameterizations. Blue: Chebyshev only; green: Chebyshev + osculating elements; red: Chebyshev + osculating elements + Fourier. Each point corresponds to a Chebyshev order (darker = higher order). Points with black outlines satisfy both error targets (3 m, 0.25 mm/s, horizontal dashed line) within 900-bits (vertical dashed line)
4.2 Message Size Computation
The message size required to broadcast the ephemeris or almanac can be estimated following the procedure outlined in [9], with additional subprocedures introduced here to compute parameter resolutions efficiently using numerical differentiation.
  1. 1.
    Resolution Computation: For each parameter (e.g., Chebyshev coefficients, Fourier coefficients, or orbital elements), determine the bit resolution required to ensure that the resulting position error remains below a specified tolerance $\delta_{x}$. The process consists of three steps:
    1. (a)
      Numerical sensitivity evaluation: Compute the sensitivity of the satellite position to each parameter by evaluating the partial derivative of the position function with respect to that parameter using finite differences:
      $\left|\frac{\partial f_{\bm{\alpha}\rightarrow x}(\bm{\alpha},t)}{\partial\alpha_{i}}\right|\approx\frac{1}{N_{t}\epsilon}\sum_{m=1}^{N_{t}}\delta f_{m}(\bm{\alpha})$ (14)
      $\delta f_{m}(\bm{\alpha})=f_{\bm{\alpha}\rightarrow x}(\bm{\alpha}+\epsilon\mathbf{e}_{i},t_{m})$
      $\qquad-f_{\bm{\alpha}\rightarrow x}(\bm{\alpha}-\epsilon\mathbf{e}_{i},t_{m})$
      where $f_{\bm{\alpha}\rightarrow x}(\bm{\alpha},t)$ maps the parameter set $\bm{\alpha}$ to the satellite position at time $t$; $t_{m}$ ($m=1,\ldots,N_{t}$) are the sampling times; $\mathbf{e}_{i}$ is the unit vector corresponding to parameter $\alpha_{i}$; and $\epsilon=10^{-3}$ is a small perturbation.
    2. (b)
      Initial bit-depth estimate: The initial estimate of the bit depth $k_{i}$ for parameter $\alpha_{i}$ is computed as
      $k_{i}=-\left\lceil\log_{2}\left(\frac{\delta_{x}}{\left|\frac{\partial f_{\bm{\alpha}\rightarrow x}}{\partial\alpha_{i}}\right|}\right)\right\rceil,$ (15)
      where $\lceil\cdot\rceil$ denotes the ceiling function and $\delta_{x}=0.01\penalty 10000\ \mathrm{m}$ is the target position accuracy.
    3. (c)
      Iterative refinement: Starting from the initial estimate in Eq. (15), iteratively adjust $k_{i}$ by $\pm 1$ until the smallest value satisfying the position accuracy constraint is found:
      $k_{i}$ $=\arg\min_{k_{i}}\bigg\{\max_{s\in\{-1,1\}}\max_{m\in[1,N_{t}]}$ (16)
      $\|f_{\bm{\alpha}\rightarrow x}(\bm{\alpha}+s2^{-k_{i}}\mathbf{e}_{i},t_{m})-\bar{\mathbf{x}}_{m}\|_{2}<\delta_{x}\bigg\}$
      where $\bar{\mathbf{x}}_{m}$ is the reference position at time $t_{m}$.
  2. 2.
    Bit Allocation: Once the required resolution is determined, the number of bits $b_{i}$ required to encode each parameter is given by
    $b_{i}$ $=\max\left\{\left\lceil\log_{2}\!\left(\alpha_{i,\mathrm{max}}-\alpha_{i,\mathrm{min}}\right)+k_{i}\right\rceil,1\right\}$ (17)
    $\quad+b_{s}+b_{m},$
    where $\alpha_{i,\mathrm{max}}$ and $\alpha_{i,\mathrm{min}}$ denote the maximum and minimum values of $\alpha_{i}$ across different fitting arcs (with varying starting mean anomalies). The terms $b_{s}$ and $b_{m}$ represent the bits allocated for sign (1 for signed, 0 for unsigned) and for margin to account for quantization and potential future adjustments, respectively. No margin bits are allocated for the bounded elements $e,i,\Omega,\omega,$ and $M$, as their ranges are limited to $[0,1]$ or $0,2\pi]$.
  3. 3.
    Total Message Size: The total message size (in bits) is the sum of the individual parameter allocations:
    $\text{Message Size}=\sum_{i=1}^{N_{c}}b_{i},$ (18)
    where $N_{c}$ is the total number of fitted parameters.
5 Ephemeris Simulation Results
5.1 Simulation Setup
The ephemeris fitting experiments are conducted for four lunar orbit regimes introduced in Section 2. We evaluate fitting intervals of 60, 120, 240, 360, and 480 min, and compare three ephemeris parameterizations: (i) Chebyshev polynomials only, (ii) Chebyshev + osculating elements, and (iii) Chebyshev + osculating elements + Fourier series.
For each orbit and fitting interval, we draw 30 starting mean anomalies uniformly over $0,2\pi)$. Using these 30 fits, we compute the required data size per representation via the bit-allocation method in Section 4.2. Fitting accuracy is reported as the 95th-percentile position (norm) and velocity (norm) errors evaluated over all points in fitting arcs with different starting mean anomalies.
Sampling for fitting uses one point per minute with Chebyshev–Robatto (Chebyshev–Lobatto) nodes; evaluation uses uniform points sampled per second.
Figure 6: Bit allocation breakdown for the 30-hour ELFO (LCRNS) using Chebyshev + osculating elements + Fourier at a 360 min fitting length.
Figure 7: Bit allocation breakdown for the 30-hour ELFO (LCRNS) using Chebyshev + osculating elements at a 360 min fitting length.
Table 2: Summary of ephemeris fitting results within a 900-bit budget (95th-percentile errors).
Orbit 120 min 240 min 360 min
Position (m) Velocity (mm/s) Bits Order Position (m) Velocity (mm/s) Bits Order Position (m) Velocity (mm/s) Bits Order
30 hr ELFO (LCRNS) 1.12E-04 1.86E-02 654 19 5.02E-03 2.68E-02 893 18 3.17E-01 6.98E-01 893 13
24 hr ELFO (LCNS) 5.37E-04 2.80E-02 816 19 5.53E-02 2.12E-01 893 17 7.63E-01 1.63E++$0$ 882 13
12 hr ELFO (LNSS) 1.30E-02 1.00E-01 882 16 2.44E-01 9.01E-01 893 15 2.28E++$0$ 5.32E++$0$ 888 12
6 hr Polar Orbit 7.41E-05 1.60E-02 579 17 8.32E-05 1.61E-02 837 29 2.06E-01 4.43E-01 889 17
5.2 Ephemeris Fitting Results for the 30-hour ELFO (LCRNS)
[…]
Consistent with prior work, augmenting Chebyshev polynomials with osculating elements yields markedly better accuracy at a fixed bit budget than Chebyshev-only fits. Moreover, for fitting ranges $\geq\!240$ min, adding a short Fourier basis to Chebyshev + elements further reduces both position and velocity errors within the 900-bit constraint. The Fourier terms efficiently capture periodic components of the perturbations, allowing lower Chebyshev orders and smaller coefficient ranges, hence fewer bits for comparable or improved accuracy. For short arcs (120 min or less), the Chebyshev + elements model often provides the best accuracy per bit, as the additional Fourier term contributes limited incremental benefit over short durations.
Across fitting lengths, the velocity constraint (0.25 mm/s, 95th) is typically more stringent than the position constraint (3 m, 95th), especially for long arcs ($\geq\!360$ min). Under the 900-bit budget, meeting the velocity target is generally feasible for fitting lengths up to 240 min.
Fig. 7 and 7 compare bit allocations at 360 min for the Chebyshev + OE + Fourier and Chebyshev + OE representations, respectively. While the Fourier model introduces roughly 120 bits for six Fourier coefficients, it enables a reduction in Chebyshev order and coefficient ranges, yielding a better overall accuracy–bit trade.
5.3 Ephemeris Fitting Results for Other Orbits
The results for the 24-hour ELFO (LCNS), 12-hour ELFO (LNSS), and 6-hour polar orbit are summarized in Table 2. Data size versus error plots for the 12-hour ELFO and the 6-hour polar orbit are shown in Figs. 8 and 9, respectively.
For ELFO regimes, trends mirror the LCRNS case: for $\geq\!240$ min, Chebyshev + elements + Fourier typically achieves the best accuracy within 900 bits; for shorter arcs, Chebyshev + elements is most efficient. For the polar orbit, Chebyshev + elements consistently outperforms the other representations across all fitting lengths, likely because short-period variations dominate and are well captured without additional Fourier terms.
As the ELFO period decreases (higher orbital speed, stronger short-period content), accuracy degrades for a fixed bit budget and fitting length, reflecting increased high-frequency variation in states and elements. Nonetheless, for all ELFOs we achieve sub-meter position and sub-mm/s velocity (95th) within 900 bits for fitting lengths up to 240 min; for the polar orbit, sub-meter and sub-mm/s accuracy remains achievable out to 360 min within the same bit budget.
Figure 8: Data size versus 95th-percentile position and velocity errors for the 12-hour ELFO (LNSS) across fitting lengths and parameterizations. Blue: Chebyshev only; green: Chebyshev + osculating elements; red: Chebyshev + osculating elements + Fourier. Points with black outlines satisfy both error targets (10 m, 2.5 mm/s, horizontal dashed line) within 900-bits (vertical dashed line)
Figure 9: Data size versus 95th-percentile position and velocity errors for the 6-hour polar orbits.
6 Almanac Fitting Results
We next evaluate the almanac model for three representative lunar orbit regimes. The primary role of the almanac is to support cold/warm starts by providing coarse satellite geometry sufficient to form a visibility list and predict approximate Doppler shifts [8]. Accordingly, we adopt a 15-day fitting interval to enable warm starts after a full lunar night (approximately 14 days).
6.1 Almanac Fitting Results for the 30-hour ELFO (LCRNS)
Fig. 10 shows the least-squares almanac fits over 15 days for the residuals of the osculating elements $\{\Delta a,\Delta e,\Delta i,\Delta\Omega,\Delta\omega,\Delta M,\Delta u\}$. The model captures the dominant secular and periodic trends, but does not fully track the highest-frequency content in the semi-major axis, which manifests as along-track (tangential) phase error—visible in the mean anomaly and argument-of-latitude traces.
Fig. 11 plots the resulting position and velocity errors over the 15-day interval. Error spikes occur near perilune, particularly in the along-track component, yet the overall errors remain within $\sim 500$ km and $50$ m/s across the interval when the satellite is visible from the lunar south pole.
Fig. 12 summarizes latitude/longitude errors of the fitted satellite orbits. Median and 99th-percentile errors remain within $1^{\circ}$ and $10^{\circ}$, respectively, over 15 days. During periods visible from the lunar south pole (elevation $5^{\circ}$), the latitude/longitude errors are typically within $3^{\circ}$, which is sufficient for discriminating visible satellites at warm start.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 10: Almanac fits for the 30-hour ELFO (LCRNS) over a 15-day interval.
Figure 11: Position and velocity errors for the 30-hour ELFO (LCRNS) almanac over 15 days. Blue: epochs visible from the lunar south pole (elevation $\geq 5^{\circ}$); red: not visible.
Figure 12: Latitude and longitude errors for the 30-hour ELFO (LCRNS) almanac over 15 days. Blue: epochs visible from the lunar south pole (elevation $\geq 5^{\circ}$); red: not visible.
6.2 Almanac Fitting Results for Other Lunar Orbits
Fig. 13 show 15-day latitude/longitude errors for the 6-hour polar orbit. Table 3 summarizes all four cases, including the 3 ELFOs.
For ELFOs, latitude/longitude errors increase as orbital period decreases (smaller semi-major axis), with LNSS showing the largest errors. Higher orbital speeds and stronger short-period content amplify high-frequency element variation, which a low-order almanac cannot fully capture. However, these larger errors arise predominantly near perilune when the satellite is often not visible from the south pole; during visible periods, the 99th-percentile latitude/longitude errors remain within $6.5^{\circ}$ across all ELFOs, which is adequate for acquisition and visible-satellite discrimination.
The polar orbit exhibits smaller and more uniform latitude/longitude errors than ELFOs. Given its near-circular geometry, the below-horizon and above-horizon error distributions are similar.
Figure 13: Latitude/longitude errors for the 6-hour polar-orbit almanac over 15 days.
Table 3: Summary of almanac errors during visibility from the lunar south pole (elevation $\geq 5^{\circ}$).
Orbit Position Error [km] Velocity Error [m/s] Latitude Error [deg] Longitude Error [deg]
Median 95% 99% Median 95% 99% Median 95% 99% Median 95% 99%
30 hr ELFO (LCRNS) 85.9 555.8 738.2 2.45 60.7 126.5 0.46 1.29 1.89 0.57 1.64 1.74
24 hr ELFO (LCNS) 169.3 594.0 783.2 4.69 101.5 161.5 0.56 1.57 2.56 0.90 1.79 2.40
12 hr ELFO (LNSS) 94.3 1006.4 1175.4 17.1 315.1 384.3 0.63 5.11 6.25 1.48 2.92 3.37
6 hr Polar Orbit 5.29 154.8 197.3 41.9 60.8 70.2 0.88 2.28 2.95 0.11 1.90 10.72
7 Conclusion
This paper presented a unified framework for designing ephemeris and almanac messages for lunar navigation satellites, addressing the unique dynamical and operational challenges of the lunar environment. By combining osculating orbital elements with Chebyshev and Fourier representations, the proposed ephemeris model achieves high-fidelity orbit reconstruction over extended fitting intervals while satisfying the LunaNet message size constraint. Across multiple orbit regimes—including 30-, 24-, and 12-hour ELFOs and a 6-hour polar orbit—the hybrid representation achieved sub-meter position and sub-millimeter-per-second velocity fitting accuracy within 900 bits for ephemeris arcs up to four hours.
For the almanac, a compact polynomial–Fourier representation of the orbital elements successfully captured long-term orbital variations over 15-day intervals, yielding median latitude and longitude errors below one degree. This level of accuracy is sufficient for distinguishing visible satellites during warm-start operations, ensuring robust acquisition performance after lunar night periods.
Overall, the proposed approach provides a scalable and interoperable foundation for generating almanac and ephemeris products for the Lunar Augmented Navigation Service (LANS). Future work will extend this framework to incorporate orbit determination and prediction errors, satellite clock models, and user receiver simulations to evaluate end-to-end navigation performance within the broader LunaNet architecture.